OpenStudy (sonyalee77):

PLEASE HELP WILL GIVE MEDAL!!!!! Graph the inequality:-5[y+4]>=3[y-4]

OpenStudy (anonymous):

it goes up very quickly. assume e is equal to 2.718281828... x is he exponent that e is raised to. equation is y = 3*e^x when x = -3, y = .15 when x = -2, y = .41 when x = -1, y = 1.1 when x = 0, y = 3 when x = 1, y = 8.15 when x = 2, y = 22.17 when x = 3, y = 60.26 when x = 4, y = 163.79 when x = 5, y = 445.24 the further positive x becomes, the further positive y = e^x becomes. the further negative x becomes, the closer to 0 y becomes. y = 3e^-20 would be equal to 6.18 * 10^-9. y = 3e^-50 would be equal to 5.78 * 10^-22 I'll send you a pic of the graph

OpenStudy (sonyalee77):

Sorry it's supposed to be on a number line but ill still give you a medal

OpenStudy (anonymous):

cool thanks (sorry I tried figuring it out)